# JEL Classification: C51, C81, D83, O13, Q12, Q54

### v1

@@ Abstract
Impact evaluations (IEs) of new technologies are critical to improving investment in national and international development goals. Yet many technologies are introduced at times or in places that lack the necessary data to conduct a well identified IE. We present a new method that combines Earth observation (EO) data, advances in machine learning, and survey data so as to allow researchers to conduct IEs when traditional economic data is missing. To demonstrate our approach, we study stress tolerant rice varieties (STRVs) introduced more than a decade ago. Using 20 years of EO data on rice production and flooding, we fail to replicate existing evidence of STRV effectiveness. We validate this failure to replicate with household panel data and through Monte Carlo simulations that demonstrate the sensitivity of past evidence to mismeasurement. Our findings speak to the challenges and promises of using EO data to conduct IEs in data poor settings.
@@ 1 Introduction
Getting oneβs hands on a sufficient amount of relevant data has always been a challenge in answering economic questions, whether theoretical or empirical. Karl Marx struggled to obtain firm-level production data for use in developing his theory of value (Marx, 1867). John Maynard Keynes lamented the lack of sufficient data to definitively disprove Arthur Pigouβs theory of unemployment (Keynes, 1936). Simon Kuznets, in part, won his Nobel Prize for creating standardized and regularly collected data on national accounts (Rockoff, 2020). One reason why the earliest advances in econometrics were made in the sub-field of agricultural economics was because George F. Warren convinced the U.S. Census Bureau to conduct a separate national census of agriculture at more frequent intervals than the decennial census (Stanton, 2007). This meant that agricultural economists had much more data for empirical research than those working on other macro or micro questions.111Nearly as important as having any data is having sufficient data. Well know empirical findings, such as the Philips Curve (Phillips, 1958) and the deterrence effect of capital punishment (Ehrlich, 1975), have proven much less robust than initially thought once additional data were added to the analysis (Hazell etΒ al., 2022; Nagin, 2013).
@@
The last quarter century has seen an explosion in the availability of data, economic and otherwise. Data, in combination with the personal computer, the internet, mobile information, and communication technology, has created a data rich world. This plethora of data has allowed economists to answer a host of new questions on anything from the historic effects of colonial institutions (Dell, 2010) to the current effects of misinformation (Allcott and Gentzkow, 2017) to the future effects of climate change (DeschΓͺne and Greenstone, 2007). Economists have also become adept at integrating traditional socioeconomic data with new types of data from a variety of sources. In particular, remotely sensed Earth observation (EO) data has become a favorite for use in economic research (Dell etΒ al., 2014; Donaldson and Storeygard, 2016). Economists have used EO data to examine questions about economic growth, agricultural productivity, land use, population growth, poverty, child mortality, and governance (BenYishay etΒ al., 2017).222A cottage industry exists of researchers using EO weather data to predict everything from human capital formation (Shah and Steinberg, 2017; Garg etΒ al., 2020) to labor markets (Chen etΒ al., 2017; Kaur, 2019; Morten, 2019) to conflict and institutions (Sarsons, 2015; KΓΆnig etΒ al., 2017) to agricultural production (Dell etΒ al., 2012; Yeh etΒ al., 2020) to intra-household bargaining power (Corno etΒ al., 2020) to technology adoption (Jagnani etΒ al., 2021; AragΓ³n etΒ al., 2021).
@@
Yet the recency of so much of this data means that there is at least one place that remains relatively data poor: the past. As Burke etΒ al. (2021) show, the increase in the spatial and temporal resolution of EO data has primarily occurred since 2010, with the most substantial improvements (m, daily) occurring in the last five years. While EO data has been used to answer many research questions about recent or contemporaneous events, its usefulness in answering economic questions about the (relatively recent) past has, thus far, been limited. The requirements for ground truth or training data to improve the accuracy of EO derived outcomes also places perceived limits on where and when EO data can be used to answer a research question.
@@
In this paper, we develop a methodology to overcome the recency bias in high quality EO data so that it can be used to answer economic research questions in environments that remain - and are likely to remain - data poor. We do this in the context of a long-term, large-scale impact evaluation on the effectiveness of stress tolerant rice varieties (STRVs) in mitigating yield loss to flooding in Bangladesh. The flood-tolerant rice variety, Swarna-Sub1, was first released in Bangladesh in 2010 and other varieties quickly followed (Mackill etΒ al., 2012). While a randomized control trial (RCT) was conducted in 128 villages in Odisha, India in 2011 (Emerick etΒ al., 2016) and household panel data on varietal-specific rice production exists in South Asia starting in 2014 (Yamano, 2017), no data exist that would allow for a before/after comparison of the impacts of adoption on a large-scale. The relatively long time since the introduction of varieties with the Sub1 gene means that the high spatial and temporal resolution EO data now being used in impact evaluations (Carlson etΒ al., 2018; Jayachandran etΒ al., 2017; Jain etΒ al., 2019) also does not exist for use as a baseline.
@@
In order to conduct our impact evaluation in this data poor environment we combine data from a variety of sources (EO, administrative, household survey), develop new approaches to generating ground truth data, and leverage recent advances in machine learning (ML). Our first task is to build maps for where rice is grown in Bangladesh. The standard approach is to collect GPS data on rice and non-rice plots and use that as training data for an ML algorithm to predict rice area at a larger scale (Burke and Lobell, 2017; Azzari etΒ al., 2021). While we have this type of ground truth data for 2021 and 2022, it does not exist for previous years. To overcome this lack of ground truth data, we hand-annotate thousands of Google Earth images to build a training data set going as far back as 2002. Having identified rice area, we next build maps of where rice was flooded. The traditional method for identifying floods is to use an existing databases of large-scale floods based on optical sensors that might be blocked by clouds (Tellman etΒ al., 2021) or to handcraft flood measures using EO data (Guiteras etΒ al., 2015; Chen etΒ al., 2017). Both methods miss many small flood events or have cloud blockage and thus under-measure the extent of flooding (Saunders etΒ al., 2023). Substantial improvements in accuracy can be achieved using ML methods similar to those used in rice mapping but here again the problem is a lack of ground truth data. To overcome this, we use radar data from Sentinel-1, a satellite with high spatial and temporal resolution data not blocked by clouds, launched in 2014, as training data for a Convolutional Neural Network - Long Short-Term Memory (CNN-LSTM) model that then predicts past flooding in lower resolution historical data (Giezendanner etΒ al., 2023; Thomas etΒ al., 2023).
@@
Our first approach to estimating the impacts of STRVs combines these rice and flood maps with the EO derived enhanced vegetative index (EVI), a common proxy for yields, and administrative data on STRV seed release. This produces a panel data set that covers the entire country for the 20-year period 2002-2021. Using a variety of econometric methods (event study, difference-in-difference, two-way fixed effects), each with its own strengths and weaknesses regarding identifying assumptions, we find little evidence that, after the introduction of STRVs, EVI in flooded areas is higher than it was in those same areas before the introduction of STRVs.
Our failure to replicate the positive impacts of STRVs on yields during floods, which have been documented in numerous agronomic trials and in an RCT, sheds doubt on the effectiveness of our EO-driven approach. To assuage this doubt, we validate our approach using household panel survey data that covers the years 2014, 2017, and 2022 and is representative of rice growing regions in Bangladesh.333Prior to the collection of the new round of survey data, we pre-specified our analysis and archived a pre-analysis plan (PAP) publicly on Open Science Foundation (OSF) (AlΒ Rafi etΒ al., 2023). We first use a two-way fixed effects (TWFE) model to estimate the impacts of household adoption of STRVs on household rice yields. That the household data all comes post-introduction of STRV means we have no baseline data in which adoption was completely exogenous (as the seed did not yet exist). Recognizing this, we implement a TWFE model with instrumental variables (IV) in which past flood events derived from our flood maps are used to predict STRV adoption. Results from our household-level analysis comport with the evidence from our EO-driven approach: there is little evidence that STRV adoption results in higher yields for the household during flooding relative to similarly severe floods experienced by the same household. Analysis using cross-sectional plot-level data corroborates the household and EO results.
Why then, using three different levels of analysis and a variety of econometric specifications, do we fail to find the large, positive impacts of STRVs on yield that exist in the agronomic and RCT data? One potential reason is that, unlike high yielding seed varieties which create unconditional yield effects, STRVs are a stochastic technology that generates higher order treatment effects. Based on agronomic field trials, the specific technology we study generates measurable treatment effects only when flooding is greater than seven days but less than 17 days. Outside that window, STRVs have yields statistically indistinguishable from non-STRVs. To test this conjecture, we conduct exploratory (not pre-specified) analysis using EO flood data on the exact number of days a region experienced flooding. We also conduct Monte Carlo simulations using the RCT data from Dar etΒ al. (2013) and Emerick etΒ al. (2016) to examine the sensitivity of those results to injecting noise into the data. This exploratory analysis demonstrates that measuring a statistically significant treatment effect for STRVs is highly sensitive to mismeasurement or mis-classification, be it in the EO or RCT data.
@@
We make three contributions to the literature. First, we contribute to the growing set of geospatial impact evaluations. A geospatial impact evaluation uses georeferenced EO data to measure either the intervention, the outcome, or both (BenYishay etΒ al., 2017). This geospatial data can be used as part of an RCT, as in Jayachandran etΒ al. (2017), (Jain etΒ al., 2019), and Huang etΒ al. (2021), but is more frequently used as part of a quasi-experimental approach to establishing the causal effects of an intervention. Geospatial impact evaluations have been used to study the impacts of governance, anti-poverty programs, and conditional cash transfers on deforestation (Burgess etΒ al., 2012; Alix-Garcia etΒ al., 2013; Ferraro and Simorangkir, 2020), and the impact of insecticide treated bednets on child mortality (Dolan etΒ al., 2019), to name a few. Many of these geospatial impact evaluations combine EO data with survey data on specific individual outcomes in specific areas. However, a growing proportion of studies, like Burgess etΒ al. (2012) and Ferraro and Simorangkir (2020), rely solely on EO and administrative data. In this paper, we combine both of these approaches, relying solely on EO and administrative data in our first set of analyses and relying on a combination of EO and household and plot survey data in the second set. We believe this multi-pronged approach provides increased reliability in our estimates by providing evidence from a variety of data sources at a variety of scales.
@@
Second, we make a methodological contribution to the broader literature that looks to combine EO data with socioeconomic data. Both the availability and quality of EO data is changing rapidly. So too is the sophistication of model building to use both optical and radar EO data to predict specific outcomes of interest. Early economic studies that integrated EO data typically relied on what Burke etΒ al. (2021) term βhandcrafted featuresβ, such as calculating ratios of reflectances at different wavelength, for use in shallow predictive models. One commonly used example of a handcrafted metric is the Modified Normalized Difference Water Index (MNDWI) (Guiteras etΒ al., 2015; Chen etΒ al., 2017; Chen and Mueller, 2018). The MNDWI is simply a ratio of the differences in surface reflectance such that positive values are inferred to be water and negative values are inferred to be land (Xu, 2006). An example of a shallow predictive model is using a simple linear regression to predict values of pixels that might be missing or obscured by clouds (Weiss etΒ al., 2020). We leverage recently developed techniques from ML to fuse optical and radar data, including models that use the spatial and temporal structure of the data, to improve predictive ability and reduce noise in our estimates of where rice is grown and where flooding occurs.
@@
Finally, we add to the literature on the impact of agricultural technology by studying a stochastic technology that generates higher order treatment effects. The aim of the first-generation of Green Revolution technologies was to improve on the low yields of landrace varieties Evenson and Gollin (2003). Newer generations of high yielding varieties now also embody genetics to allow the plant to tolerate biotic and abiotic stresses (Hossain etΒ al., 2006). This creates a challenge for the researcher trying to measure impact, as the effectiveness of STRVs manifests only within a very specific window and whether an observations fits within that window is often a function of a stochastic event. As our results show, this βGoldilocks Problemβ places a limit on the effectiveness of geospatial impact evaluations, particularly in data poor settings. However, as our exploratory analysis and Monte Carlo simulations show, this sensitivity to misallocation, mismeasurement, or mis-classification is not limited to EO data but applies to survey and experimental data as well - a point made by Carletto etΒ al. (2015), Abay etΒ al. (2019), and Alix-Garcia and Millimet (2033) that our paper reinforces.
@@ 2 Study Context
Green Revolution technologies played a critical role in improving food security and reducing poverty in many developing countries (Evenson and Gollin, 2003). Early in the Green Revolution, many technologies were developed for and disseminated in agronomically favorable environments, bypassing environments where abiotic stresses resulted in low or uncertain yields (Khush and Virk, 2000; Gollin etΒ al., 2021). These earlier technologies often embodied unconditional yield effects. Simply planting them resulted in high yields under existing growing conditions and farming practices, though even greater yields could be achieved when high yielding seeds were bundled with other modern inputs.
@@
Stress tolerant seed varieties, including stress tolerant rice varieties (STRVs), were developed to address the rising challenges of climate change-induced stresses, such as salinity, extreme temperature, drought, or submergence (Lichtenthaler, 1998; Sevanthi etΒ al., 2019). In South and South-East Asia, where seasonal flooding is critical to rice cultivation, farmers face a substantial risk of crop loss due to early and sustained flooding (Mishra etΒ al., 2015). To address this, researchers isolated submergence tolerant traits (the Sub1 gene) in wild rice (Xu and Mackill, 1996) to create Swarna-Sub1 (Yamauchi etΒ al., 2013).
@@
Swarna-Sub1 has no yield penalty in the absence of flooding. If and when flooding occurs, Swarna-Sub1 can survive complete submergence of between seven and 17 days, in up to 25 cm of water, with little yield penalty. But, while Swarna and other Sub1 varieties outperform non-STRVs, the yield effect is not unconditional. FigureΒ 1 traces the yield effect in relation to the duration of flooding. There is a clear βsweet spotβ in which Sub1 and other STRVs produce large yield gains. The effectiveness of STRVs is a function of an essentially random event (the length and intensity of flooding) making STRVs a stochastic technology and meaning attempts to quantify their impact requires measuring a higher order treatment effect. If flooding is outside the 10 day window (less than seven or more than 17 days), then we would expect STRVs to have a treatment effect equal to zero.444Additional details on Sub1 varieties is available in Online AppendixΒ A.

### v2

@@ Abstract
New technologies are sometimes introduced at times or in places that lack the necessary data to conduct a well-identified impact evaluation. We develop a methodology that combines Earth observation (EO) data and advances in machine learning with administrative and survey data so as to allow researchers to conduct impact evaluations when traditional economic data is missing. To demonstrate our method, we study stress tolerant rice varieties (STRVs) first introduced to Bangladesh 15 years ago. Using EO data on rice production and flooding for the entire country, spanning two decades, we find evidence of STRV effectiveness. We highlight how the nature of the technology, which is only effective under a specific set of circumstances, creates a Goldilocks Problem that EO data is particularly well suited to addressing. Our findings speak to the promises and challenges of using EO data to conduct impact evaluations in data poor settings.
@@ 1 Introduction
Obtaining a sufficient amount of relevant data has always been a challenge in answering economic questions, whether theoretical or empirical. Karl Marx struggled to obtain firm-level production data for use in developing his theory of value (Marx, 1867). John Maynard Keynes lamented the lack of sufficient data to definitively disprove Arthur Pigouβs theory of unemployment (Keynes, 1936). Simon Kuznets, in part, won his Nobel Prize for creating standardized and regularly collected data on national accounts (Rockoff, 2020). One reason why the earliest advances in econometrics were made in the field of agricultural economics was because George F. Warren convinced the U.S. Census Bureau to conduct a separate national census of agriculture at more frequent intervals than the decennial census (Stanton, 2007). This meant that agricultural economists had much more data for empirical research than those working on macro or other micro questions.111Nearly as important as having any data is having sufficient data. Well know empirical findings, such as the Philips Curve (Phillips, 1958) and the deterrence effect of capital punishment (Ehrlich, 1975), have proven much less robust than initially thought once additional data were added to the analysis (Hazell etΒ al., 2022; Nagin, 2013).
@@
The last quarter century has seen an explosion in the availability of data, both economic and otherwise. Data, in combination with the personal computer, the internet, mobile information, and communication technology, has created a data rich world. This plethora of data has allowed economists to answer a host of new questions on anything from the historic effects of colonial institutions (Dell, 2010) to the current effects of misinformation (Allcott and Gentzkow, 2017) to the future effects of climate change (DeschΓͺne and Greenstone, 2007). Economists have also become adept at integrating traditional socioeconomic data with new types of data from a variety of sources. In particular, remotely sensed Earth observation (EO) data has become a favorite for use in economic research (Dell etΒ al., 2014; Donaldson and Storeygard, 2016). Economists have used EO data to examine questions about economic growth, agricultural productivity, land use, population growth, poverty, child mortality, and governance (BenYishay etΒ al., 2017).
@@
Yet the recency of so much of this data means that there is at least one place that remains relatively data poor: the past. As Burke etΒ al. (2021) show, the increase in the spatial and temporal resolution of EO data has primarily occurred since 2010, with the most substantial improvements (m, daily) occurring in the last decade. While EO data has been used to answer many research questions about recent or contemporaneous events, its usefulness in answering economic questions about the (relatively recent) past has, thus far, been limited. The requirements for ground truth or training data to improve the accuracy of EO derived outcomes also places perceived limits on where and when EO data can be used to answer a research question.
@@
In this paper, we develop a methodology to overcome the recency bias in high quality EO data so that it can be used to answer economic research questions in settings that remain - and are likely to remain - data poor. We do this in the context of a long-term, large-scale impact evaluation on the effectiveness of stress tolerant rice varieties (STRVs) in mitigating yield loss to flooding in Bangladesh. The flood-tolerant rice variety, Swarna-Sub1, was first released in Bangladesh in 2010 and other varieties quickly followed (Mackill etΒ al., 2012). While a randomized control trial (RCT) was conducted in 128 villages in Odisha, India in 2011 (Emerick etΒ al., 2016) and household panel data on varietal-specific rice production exists in South Asia starting in 2014 (Yamano, 2017), no data exist that would allow for a before/after comparison of the impacts of adoption on a large-scale. The relatively long time since the introduction of varieties with the Sub1 gene means that the high spatial and temporal resolution EO data now being used in impact evaluations (Carlson etΒ al., 2018; Jayachandran etΒ al., 2017; Jain etΒ al., 2019) also does not exist for use as a baseline.
Further impoverishing the datasetting is the fact that STRVs, like many new agricultural technologies, have a stochastic treatment effect. For the technology to work and, by extension, be identifiable from space, it requires not too much flooding but also not too little flooding. This creates a Goldilocks Problem in which the treatment effect of the technology is only be observable under a specific set of circumstances. For varieties with the Sub1 gene, that set of circumstances is complete submergence for longer than five days but less than 20. Outside that 5-20 day window, STRVs have yields statistically indistinguishable from non-STRVs. Thus, in any existing data for use in an impact evaluation, one would need precise information about the location and duration of flooding. Given the now well-known problem of measurement error in self-reported agricultural data (Wollburg etΒ al., 2021; Abay etΒ al., 2023), households may be unable to retrospectively report on flooding with the precision necessary to capture the agronomically relevant level of inundation required for STRVs to be effective. As we show using Monte Carlo simulations and agronomic field trial data, identifying these higher order treatment effects is highly sensitive to classical measurement error.
@@
In order to conduct our impact evaluation in this data poor environment we combine data from a variety of sources (EO, administrative, household survey) and leverage recent advances in machine learning (ML). Our first task is to build maps for where rice is grown in Bangladesh. The standard approach is to collect GPS data on rice and non-rice plots and use that as training data for an ML algorithm to predict rice area at a larger scale (Burke and Lobell, 2017; Azzari etΒ al., 2021). While we have this type of ground truth data for 2021 and 2022, it does not exist for previous years. To overcome this lack of ground truth data, we hand-annotate thousands of Google Earth images to build a training dataset going as far back as 2002. Having identified rice area, we next build maps of where rice was flooded. The traditional method for identifying floods is to use an existing databases of large-scale floods based on optical sensors that might be blocked by clouds (Tellman etΒ al., 2021) or to handcraft flood measures using EO data (Guiteras etΒ al., 2015; Chen etΒ al., 2017). Both methods miss many small flood events or have cloud blockage and thus under-measure the extent of flooding (Saunders etΒ al., 2023). Substantial improvements in accuracy can be achieved using ML methods similar to those used in rice mapping but here again the problem is a lack of ground truth data. To overcome this, we use radar data from Sentinel-1, a satellite with high spatial and temporal resolution not blocked by clouds, launched in 2014, as training data for a Convolutional Neural Network - Long Short-Term Memory (CNN-LSTM) model that then predicts past flooding in lower spatial resolution historical data using the MODIS sensor and topographic variables (Giezendanner etΒ al., 2023; Thomas etΒ al., 2023).
@@
Our primary approach to estimating the impacts of STRVs combines these rice and flood maps with the EO derived enhanced vegetative index (EVI), a common proxy for yields, and administrative data on STRV seed release. This produces a panel dataset that covers the entire country for the period 2002 through 2021. We use a variety of econometric methods (event study, difference-in-difference, two-way fixed effects), each with its own strengths and weaknesses regarding identifying assumptions. In order to highlight the severity of the Goldilocks Problem in our setting, we vary the level of inundation required to be considered a flood. And we vary the length and duration of the window in which a flood occurs. This inherently agnostic approach to what constitutes a flood produces a large number of flood metrics that are potentially βjust rightβ for identifying the treatment effects of STRVs. While a majority of these flood metrics are too little or too much flooding to identify significant impacts, we do find a set of flood metrics that produce positive and significant results. These Goldilocks floods produce consistent results across variation in how we measure EVI, our model specification, and if we use households panel data, though not to different spatial aggregations of the data.
@@
We make three contributions to the literature. First, we contribute to the growing set of geospatial impact evaluations. A geospatial impact evaluation uses georeferenced EO data to measure either the intervention, the outcome, or both (BenYishay etΒ al., 2017). This geospatial data can be used as part of an RCT, as in Jayachandran etΒ al. (2017), Jain etΒ al. (2019), and Huang etΒ al. (2021), but is more frequently used as part of a quasi-experimental approach to establishing the causal effects of an intervention. Geospatial impact evaluations have been used to study the impacts of governance (Burgess etΒ al., 2012), anti-poverty programs (Alix-Garcia etΒ al., 2013), conditional cash transfers (Ferraro and Simorangkir, 2020), and the impact of insecticide treated bed nets (Dolan etΒ al., 2019), to name a few. Many of these geospatial impact evaluations combine EO data with survey data on specific individual outcomes in specific areas. However, a growing proportion of studies, like Burgess etΒ al. (2012) and Ferraro and Simorangkir (2020), rely solely on EO and administrative data. In this paper, our primary results relying solely on EO and administrative data. We then show that these results are robust to using a combination of EO and household survey data. We believe this multi-pronged approach provides increased reliability in our estimates by providing evidence from a variety of data sources at a variety of scales.
@@
Second, we make a methodological contribution to the broader literature that looks to combine EO data with socioeconomic data. Both the availability and quality of EO data is changing rapidly. So too is the sophistication of model building that uses both optical and radar EO data to predict specific outcomes of interest. Early economic studies that integrated EO data typically relied on what Burke etΒ al. (2021) term βhandcrafted featuresβ for use in shallow predictive models. One commonly used example of a handcrafted metric is the Modified Normalized Difference Water Index (MNDWI) (Guiteras etΒ al., 2015; Chen etΒ al., 2017; Chen and Mueller, 2018). The MNDWI is simply a ratio of the differences in surface reflectance such that positive values are inferred to be water and negative values are inferred to be land (Xu, 2006). An example of a shallow predictive model is using a simple linear regression to predict values of pixels that might be missing or obscured by clouds (Weiss etΒ al., 2020). We leverage recently developed techniques from ML to fuse optical and radar data, including models that use the spatial and temporal structure of the data, to improve predictive ability and reduce noise in our estimates of where rice is grown and where flooding occurs.
@@
Finally, we add to the literature on the impact of agricultural technology by studying a stochastic technology that generates higher order treatment effects. The aim of the first-generation of Green Revolution technologies was to improve on the low yields of landrace varieties (Evenson and Gollin, 2003). Newer generations of high yielding varieties now also embody genetics to allow the plant to tolerate biotic and abiotic stresses (Hossain etΒ al., 2006). This creates a challenge for the researcher trying to measure impact, as the effectiveness of these technologies only manifests within a very specific set of circumstances that is often a function of a stochastic event. Identifying the treatment effects of these stochastic technologies therefore is highly sensitivity to mismeasurement, misallocation, or mis-classification. The existence of measurement error in both self-reported data (Wollburg etΒ al., 2021; Abay etΒ al., 2023) and in EO data (Alix-Garcia and Millimet, 2023; Josephson etΒ al., 2025), means that there are many settings in which data might exist but not at a sufficient level of precision to be useful for capturing higher order treatment effects. The Goldilocks Problem created by stochastic technologies expands the scope and scale of data poor settings. Our paper provides new tools to improve the precision of existing data so as to allow researchers to answer these type of questions.
@@ 2.1 From the Green Revolution to the Goldilocks Problem
Green Revolution technologies played a critical role in improving food security and reducing poverty in many developing countries (Evenson and Gollin, 2003). Early in the Green Revolution, many technologies were developed for and disseminated in agronomically favorable environments, bypassing environments where biotic and abiotic stresses resulted in low or uncertain yields (Khush and Virk, 2000; Gollin etΒ al., 2021). These earlier technologies often embodied unconditional yield effects. Simply planting them resulted in higher yields under existing growing conditions and farming practices, though even greater yields could be achieved when high yielding seeds were bundled with other modern inputs.
@@
Stress tolerant seed varieties, including stress tolerant rice varieties (STRVs), were developed to address the rising challenges of climate change-induced stresses, such as salinity, extreme temperature, drought, or submergence (Lichtenthaler, 1998; Sevanthi etΒ al., 2019). In South and Southeast Asia, where seasonal flooding is critical to rice cultivation, farmers face a substantial risk of crop loss due to early and sustained flooding (Mishra etΒ al., 2015). To address this, researchers isolated submergence tolerant traits (the Sub1 gene) in wild rice (Xu and Mackill, 1996) and cross-bred it with a high yielding variety (Swarna) to create Swarna-Sub1 (Yamauchi etΒ al., 2013).
@@
Swarna-Sub1 has no yield penalty in the absence of flooding. If and when flooding occurs, Swarna-Sub1 can survive complete submergence anywhere from between five and 20 days, in up to 25 cm of water, with little yield penalty. But, while Swarna-Sub1 and other Sub1 varieties outperform non-STRVs, the yield effect is not unconditional. FigureΒ 1 traces the yield effect in relation to the duration of flooding. There is a clear region that can be considered the βjust rightβ amount of flooding for Sub1 and other STRVs to produce large yield gains. But what exactly is βjust rightβ flooding varies in both the start and end day of the flood window (length) as well as the days within that window (duration). Some studies report the window as seven to 15, others as five to 17, and still others five to 20. This means there is disagreement about not only when the window starts (five, seven, etc.) and ends (15, 20, etc.) but the number of days that make up that window. Thus, the effectiveness of STRVs is a function of an essentially random event (the length and duration of flooding) making STRVs a stochastic technology and meaning attempts to quantify their impact requires measuring a higher order treatment effect. If flooding is outside some window of day that varies between five and 20, then we would expect STRVs to have a treatment effect equal to zero.222Additional details on Sub1 varieties is available in Online AppendixΒ A.
2.2 Assessing Sensitivity
The manifestation of the Goldilocks Problem for STRVs means that there is variation around what flooding may be βjust right.β Yields will vary based on the duration of submergence but also the start and end of the flood window. Beyond the length of time in the flood window and the size of the flood window, there are issues regarding the timing of floods in the growing season and the intensity or depth of flooding. This variation around what exactly characterizes a Goldilocks flood means that the researcher has a large set of potential ways to define a flood.333To avoid this problem, we pre-specified our analysis and archived a pre-analysis plan on Open Science Foundation (OSF) (AlΒ Rafi etΒ al., 2023). Results from the original, pre-specified approach are publicly available in our populated pre-analysis plan (Michler etΒ al., 2024). None of the results in this paper were pre-specified, meaning all results can be considered exploratory. Information on our rationale for the pre-analysis plan as well as its implementation can be found in Online AppendixΒ B.
To demonstrate the degree to which measuring STRV impact is sensitive to capturing just the right amount of flooding (and to mismeasurement in general), we conduct a series of Monte Carlo simulations using experimental data that forms the basis of Dar etΒ al. (2013) and Emerick etΒ al. (2016).444Though the two papers rely on the same source of data, the regressions run in Dar etΒ al. (2013), which closely match tests for efficacy in the agronomic field trial literature, are not the specifications used in Emerick etΒ al. (2016). We are not replicating any specific result in Emerick etΒ al. (2016). Rather we are using the full two years of data to reproduce results that match Dar etΒ al. (2013) in magnitude, sign, and significance. We start by calculating the mean and standard deviation of flood duration in the RCT data (; ) . Next, we draw, with replacement, from a normal distribution with mean zero and standard deviation equal to one percent of the standard deviation in the data . For each observation of flood duration, we add one of these randomly drawn values and then estimate the impact of STRVs during flooding by regressing the yield data on indicators for variety type, the new duration of flooding, and the interaction of variety with the new duration of flooding. We repeat this process 10,000 times. We then increase the noise by one percent, so that the new distribution that we draw from has mean zero and a standard deviation equal to two percent of the original standard deviation . We add this new noise level to the original data and rerun the regression for 10,000 draws from this new distribution. We conduct this process at one percent intervals from up to , meaning we run 2,000,000 simulations.
We conduct the Monte Carlo simulations first for flood duration and then, independently, for yield. Results for both are presented in the top two panels of FigureΒ 2, which graph the distribution of -values on the interaction of STRV adoption and flooding for each level of added noise. The vertical dashed red line marks . In the top row for both panels, all -values from the 10,000 simulations are statistically significant. This is because we are just replicating Dar etΒ al. (2013) 10,000 times. When we add to the true flood data, of -values are still significant. But adding just two percent of the true standard deviation as noise to the flood data results in only of regressions producing significant results. Adding any more than six percent of the true standard deviation to the true flood data results in less than of -values being significant - the same amount of -values one would expect to be significant due to random chance with a critical value. Results are similar for yield, but the loss of significance happens slightly slower, with of -values significant after adding to the flood data.
We follow a similar process for simulating mis-classification in the adoption data. However, as adoption is a binary indicator, it is not useful to use the value of the standard deviation to add noise to the data. Rather, we calculate the percentage of adopters in the data and then re-assign adopters and non-adopters at random. So, for one percent of the data we draw at random four households, determine if they are an adopter or non-adopter, and then switch their adoption status. We repeat this reassignment times. As before, we start the process at zero, so that we replicate the original results, and then move upwards at one percent intervals until we are re-assigning half of all observations.
Results from the mis-classification simulations are graphed in the bottom panel of FigureΒ 2. It is clear that the results are much more robust to mis-classification of adoption status. Mis-classifying up to five percent of observations has no meaningful impact on the outcomes. Mis-classifying of outcomes results in about half of all regressions producing significant results. Only if one mis-classified of observations does the share of significant -values fall below the ten percent threshold, where their significance could be wholly ascribed to random chance.
The take-away from this exercise is that even results on the efficacy of STRVs using experimental data can be very sensitive to mis-measurement (though less so to mis-classification). Perturbing the true flood data by adding to each value a random number drawn from a distribution , or two percent of the standard deviation, produces non-significant results in of regressions. Significance goes away adding slightly more noise to the yield data. In terms of what this means for our study context, the Monte Carlo simulations demonstrate the need for not only sufficient data, but sufficiently precise data, if one is to identify the stochastic treatment effects of STRVs. Not only do we need data to establish a pre-2010 baseline for our study, but we need that data to contain precise information on flooding, something Dar etΒ al. (2013) and Emerick etΒ al. (2016) were able to collect in a small-scale experimental setting, but which is rare in nationally representative observational data or census data. Only through highly refined EO data could we ever hope to measure the relevant flooding that occurred more than 20 years ago.
2.3 A Data Poor Setting